Seth Bullock, the owner of Bullock Gold Mining, is evaluating a new gold mine in South Dakota. Dan Dority, the company’s geologist, has just finished his analysis of the mine site. He has estimated that the mine would be productive for eight years, after which the gold would be completely mined. Dan has taken an estimate of the gold deposits to Alma Garrett, the company’s financial officer. Alma has been asked by Seth to perform an analysis of the new mine and present her recommendation on whether the company should open the new mine. Alma has used the estimates provided by Dan to determine the revenues that could be expected from the mine. She has also projected the expense of opening the mine and the annual operating expenses. If the company opens the mine, it will cost $ 635million today, and it will have a cash outflow of $ 45 million nine years from today in costs associated with closing the mine and reclaiming the area surrounding it. The expected cash flows each year from the mine are shown in the table. Bullock Mining has a required return of 12 percent on all of its gold mines. Year Cash Flow 0 − $ 635,000,000 1 89,000,000 2 105,000,000 3 130,000,000 4 173,000,000 5 205,000,000 6 155,000,000 7 145,000,000 8 122,000,000 9 − 45,000,000 QUESTIONS 1. Construct a spreadsheet to calculate the payback period, internal rate of return, modified internal rate of return, and net present value of the proposed mine.
Answer #1

Note: Amounts are in $.

Years Cash Flow Cumulative Cash Flow Payback Period Formula for the column to the left
0 -635000000 -635000000
1 89000000 -546000000 Not yet paid back =IF(AND(C2<0,C3>=0),A2+ABS(C2/B3),"Not yet paid back")
2 105000000 -441000000 Not yet paid back =IF(AND(C3<0,C4>=0),A3+ABS(C3/B4),"Not yet paid back")
3 130000000 -311000000 Not yet paid back =IF(AND(C4<0,C5>=0),A4+ABS(C4/B5),"Not yet paid back")
4 173000000 -138000000 Not yet paid back =IF(AND(C5<0,C6>=0),A5+ABS(C5/B6),"Not yet paid back")
5 205000000 67000000 4.673170732 =IF(AND(C6<0,C7>=0),A6+ABS(C6/B7),"Not yet paid back")
6 155000000 222000000
7 145000000 367000000
8 122000000 489000000
9 -45000000 444000000
Total 444000000
IRR 13.02% Formula =IRR(B2:B11)
MIRR 12.48% Formula =MIRR(B2:B11,13.02%,12%)
NPV @ 12% 21547295.32 Formula =NPV(12%,B2:B11)
Payback Period 4.67 Years (as calculated above)

Hope this helps. feel free to ask for clarifications. Kindly consider leaving a thumbs up - that will enable us to keep solving.